2x^2 + 7x-9=2x^2-18
Answer by saint(2)

2x^2 + 7x-9=2x^2-18
equate to zero
2x^2 + 7x-9-2x^2+18=0
7x-9+18=0
7x+9=0
7x=-9
x=-9/7